### v1

@@ 3.4 Deriving the super-critical manifold
(22) $\mathcal{Z}=\left\{(u,v,\theta_{0})\in\mathbb{R}^{2}\times[0,2\pi)\,|\,v=F(u),\ u-bv+\mu-Eb\cos(\theta_{0}-\pi/2)=0\right\}.$
@@
(23a) 00 0 $=-v+F(u)$
(23b) 00 0 $=u-bv+\mu-Eb\cos\left(\theta-\pi/2\right)$
(23c) $\frac{d\theta}{d\tau_{2}}$ $=1$
@@
(24) $J_{SL}=\begin{bmatrix}\varepsilon^{-1}F^{\prime}(u)&-\varepsilon^{-1}\\ 1&-b\end{bmatrix}$
@@
Converting to the original $(x,y,\theta)$ coordinates (with $x=u-1$, see Sec. 3.1), a Hopf bifurcation occurs at $x_{dh,\pm}=\pm\sqrt{1-\varepsilon b}$ along $\mathcal{Z}$. This Hopf bifurcation does not belong to the original system (2), but instead to the slow layer problem (21). Such bifurcations are called delayed Hopf bifurcations [27].
Remark 3.10.
@@ 4 Spike-adding mechanisms in the FHN model with periodic forcing
4.1 Region of interest for bursting behavior in the $(\omega,E)$-plane
First, we compare our theoretical results (summarized in Figure 3) with our numerical observations of the FHN model with periodic input (shown in Figure 2). We find that parametric regions where bursting solutions maintain a given number of spikes, and the boundaries between them, appear primarily between the curves of folded saddle-nodes on $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$ respectively; see Figure 5. This area is the union of regions II and III where the model has two and only two folded singularities (saddle plus node/focus), with both of them located on $\mathcal{L}_{-}$.
An exemplar bursting trajectory of the FHN model (1), projected on the $(x,y,\theta)$ space, is illustrated in Figure 4. The figure was generated by simulating system (1) for a pair of parameters $(\omega,E)$ chosen in region II. Note that $x$, $y$ in (1) are calculated from $u$, $v$ and $\theta$ in (3) as described in Sec. 3.1.
@@ 4.1 Region of interest for bursting behavior in the $(\omega,E)$-plane
To conclude, we will work in the remaining sections of the paper (Sec. 4.3 and beyond) under the assumption that $E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}$ see Figure 5; gray area, including blue-colored regions.
Figure 5: Parameter region of interest for bursting. Cusp-like boundaries for changing spike counts are located where a folded node and folded saddle exist on $\mathcal{L}_{-}$ (region II in Figure 3; see also Figure 2). The regions below the saddle-node (SN) bifurcation on $\mathcal{L}_{-}$ ($E=E^{*}_{\ell,\delta}$) and above the curve marking the transition from folded node to folded focus ($E=E^{**}_{\ell,\delta}$) are grayed-out. The parameter region above the SN on $\mathcal{L}_{+}$ ($E=E^{*}_{r,\delta}$) is shown in white. In our analysis we constrained input forcing parameters to $E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}$. Recall that $\omega=\varepsilon\delta$ with $\varepsilon=0.08$.
@@ 4.2 Complementary FHN dynamics in other input-based regions of the parameter plane
Before analyzing bursting patterns in the FHN model with periodic forcing, we review again the results obtained in Figure 3. The figure illustrates six distinct regions of the $(\omega,E)$ parameter plane, characterized by the stability types of folded equilibria and their existence when crossing $\mathcal{L}_{-}$ or $\mathcal{L}_{+}$.
Region I contains the parameters for which the system admits no folded equilibria, and the system admits no bursting solutions.
@@ 4.4 Adding spikes through the folded saddle canard
By Proposition 3.2, the desingularized system (10) admits a folded saddle on $\mathcal{L}_{-}$ as long as $E>E^{*}_{\ell,\delta}$. However, the desingularized system was obtained by a local state-dependent time rescaling that reverses the direction of the flow on $\mathcal{M}_{r}$ compared to the flow of (9) and (1). As such, desingularized solutions beginning in $\mathcal{M}_{r}$ that flow towards the saddle point $P_{S\ell}$ of coordinates $(u_{S\ell},v_{S\ell},\theta_{S\ell})$ (see Proposition 3.2) are reversed in the slow-reduced and in the full problem. In particular, the points that lie on $W_{S\ell}^{s}\cap\mathcal{M}_{r}$, where $W_{S\ell}^{s}$ is the stable manifold of the saddle, are repelled in the full system (1) away from $P_{S\ell}$. The curve $W_{S\ell}^{s}$ provides a path for slow solutions to pass through the saddle from $\mathcal{M}_{a-}$ into $\mathcal{M}_{r}$. This curve forms a vrai canard (or true canard), which perturbs to a canard segment in $\mathcal{M}_{a-}^{\varepsilon}\cap\mathcal{M}_{r}^{\varepsilon}$ for $\varepsilon>0$ [35].
For the desingularized system (10) the vrai canard acts as a separatrix, distinguishing solutions that pass between $\mathcal{M}_{a-}$ and $\mathcal{M}_{r}$ from those that do not. For the full system (1) the canard serves as a threshold, distinguishing solutions that remain in the neighborhood of $P_{S\ell}$ while crossing into $\mathcal{M}_{r}^{\varepsilon}$ from those that jump across to reach $\mathcal{L}_{+}$ immediately [41, 40].
Figure 7: Spike adding at folded saddle. Trajectories from the full three-dimensional system (3) projected onto the $\theta$ – $x$ plane for $E=0.482$ and (A) $\omega=0.02206875$, (B) $\omega=0.0220625$, and (C) $\omega=0.02$ (all in Region II of Figure 3). As a trajectory nears the termination of a burst sequence, its intersection with $\mathcal{M}_{a-}$ approaches $W^{s}_{S\ell}$. The proximity to $W^{s}_{S\ell}$ determines if there is a jump-back canard (A, no spike), a jump across canard (B, spike), or a jump at the fold curve (C, spike) followed by a subsequent return to the super-slow manifold $Z$ (in purple). The transition from jump-back to jump-across canard marks the addition of a spike near the folded saddle (indicated by a circle). In panel B (resp. C) the trajectory from A (resp. B) is shown in gray for comparison. In the desingularized system, flows are reversed along $\mathcal{M}_{r}$; the stable manifold of the saddle $W^{s}_{S\ell}$ (in blue) and the unstable manifold of the saddle $W^{u}_{S\ell}$ (in red; very close to $Z$) are drawn as dashed lines in this region.
@@
If no member of the $\theta$-sequence is close to the intersection of the separatrix with $\Sigma_{S\ell,-2}$, then we do not consider spike-adding to occur at the saddle. In this case, upon reaching $\theta^{m-1}_{-2}$, a solution flows towards the fold line $\mathcal{L}_{-}$ and makes a fast jump (relaxation oscillation) at it. The return dynamics to $\mathcal{M}_{a-}^{\varepsilon}$ bring the solution well past separatrix, leading it back to $\mathcal{Z}^{\delta}$, as in the previous cases (Figure 7C).
@@ 4.5 Adding spikes through the folded node canard
At the onset of an input period, bursting solutions are funneled towards the repelling region of the slow manifold through the folded node $P_{N\ell}$ of coordinates $(u_{N\ell},v_{N\ell},\theta_{N\ell})$ (see Proposition 3.2). Upon exiting the node, trajectories remain near the repelling manifold, forming canard segments at the start of a burst. The segment escapes the repelling region by either jumping back to the lower attracting branch where it originated (Figure 8A, B; see the trajectory immediately past the square mark), or by jumping across to the upper attracting branch (Figure 8C):
@@
A jump-back out of the folded node does not necessarily decrease the total number of spikes possible in a bursting oscillation. Suppose, for example, that an initial jump-across occurs and $n-1$ spikes follow (giving a total of $n$ spikes) before spike-termination at the saddle. A fast jump-back and subsequent return to the fold curve may leave enough time (or phase $\theta(t)$) for $n$ spikes before termination at the saddle. To alter the number of spikes in a burst, the initial jump-back must significantly change the trajectory’s ultimate proximity to $W_{S\ell}^{s}$, the stable manifold of the saddle, to undergo the spike-changes described in the previous section.
@@
Figure 8: Spike adding at folded node. Trajectories from the full three-dimensional system (3) projected onto the $\theta$ – $x$ plane for $E=0.482$ and (A) $\omega=0.0236$, (B) $\omega=0.02506875$, and (C) $\omega=0.025075$ (all in Region II of Figure 3). Trajectories are funneled towards the folded node (square) and into $\mathcal{M}_{r}$. After an excursion along $\mathcal{M}_{r}$, trajectories either jump back towards $\mathcal{M}_{a-}$ (A, B, no spike) or jump across towards $\mathcal{M}_{a+}$ (C, spike). From A to B, the escape from $\mathcal{M}_{r}$ towards $\mathcal{M}_{a-}$ yields a small amplitude oscillation with increasing magnitude. From B to C, the small amplitude oscillation gives rise to a jump-across towards $\mathcal{M}_{a+}$. The jump-across marks the addition of a spike out of the folded node canard. In panel B (resp. C) the trajectory from A (resp. B) is shown in gray for comparison. In the desingularized system, flows are reversed along $\mathcal{M}_{r}$, so the stable manifold of the saddle (circle) which is $W^{s}_{S\ell}$ (in blue), and the unstable manifold $W^{u}_{S\ell}$ of the saddle (in red; very close to the super-critical manifold $Z$) are drawn as dashed lines in this region.
Remark 4.1.
Folded node canard dynamics occur exclusively in Region II of Figure 3. As the parameter regime transitions from Region II into III, the folded node becomes a folded focus and the canards exist only at the folded saddle. The cusps in the spike-count boundaries appear only in Region II as well (see Figure 5) -- they arise from the complex interaction of the jump-back dynamics at the folded node with the jump-across dynamics at the folded saddle. In Region III where there is a folded focus, the spike-adding is influenced by the decay rate of the stable spirals near the focus and the saddle-canard dynamics. However, these interactions do not induce cusp-like boundaries (transparent-blue area in Figure 5). In the next section we examine the interaction between the folded node and folded saddle dynamics in more detail.
@@ 5.1 Level sets of trajectory $L^{2}$ norm
(25) $\left(\frac{1}{T}\int_{0}^{\top}\left[x(t)^{2}+y(t)^{2}\right]\,\mathrm{d}t\right)^{1/2}$
@@
These two organizing features -- the asymptotic curves and the cusp-connecting level sets -- depict how the cusps form in the spike-count boundaries. Fundamentally, cusps occur in parameter regimes in which the bifurcation between the jump-back and jump-across canards at the node leads to trajectories that ultimately land close enough to the stable manifold of the saddle to be influenced by saddle-canard dynamics. Put another way, the cusps arise as the bifurcation curve for spike-adding at the node (Figure 9; see the curves depicted by the blue arrows) collides with the bifurcation curve for spike-adding at the saddle (Figure 9; see the curves depicted by red arrows). Thus, we attribute the cusp formation to the interaction of the folded node and folded saddle dynamics.
@@ 5.2 Phase distance
The mechanisms of spike-adding at the folded node and folded saddle indicate how the spike-count boundaries might be derived analytically. Estimating the time spent (or ‘phase’ spent when considering the $\theta$ – $x$ plane) on $\mathcal{M}_{a-}^{\varepsilon}$ before the first jump on the fast timescale gives an approximation for the delay-to-spike within a burst. Similarly, estimating where the stable manifold of the saddle intersects the line $x=-2$ allows an approximation for spike termination (recall that $\Sigma_{S\ell,-2}$ denoted the plane containing this point). In particular, one would consider the phase difference between these two estimates, relative to the intra-burst spiking frequency, to estimate the number of spikes in a burst.
Figure 10: Estimating duration of spiking phase. A sample trajectory’s return to the line $x=-2$ following its first spike is marked as $\theta^{1}_{-2}$. The intersection of the saddle’s stable manifold with $x=-2$ is marked as $\theta_{S\ell,-2}$. The difference between these values, $\Delta\theta$, was used to estimate the number of spikes in a burst; see Eq. (26). Circle/square markers and the color code for different curves are the same as in previous figures.
This approach is hindered by a few factors. First, predicting the bifurcation point between a jump-back and a jump-across canard at the node is difficult, as these dynamics likely involve bifurcations of canard solutions [6, 39]. Second, the stable manifold of the saddle comes from the singular limit as $\varepsilon\to 0$, making it hard to exactly predict the separatrix that distinguishes jump-back from jump-across solutions at the saddle. And third, we do not establish an analytic relationship between the intra-burst spiking frequency and input parameters.
@@
(26) number of $\text{number of spikes}=1+\left[\frac{\Delta\theta}{1000\times\omega}\right]\times f_{burst}\,.$
For numerical calculations, we set $f_{burst}=27\,\mathrm{Hz}$ as this was the predominant frequency observed in the constantly forced system. Within formula (26), we round to the next largest integer as a fractional quantity still allows for an additional spike at the saddle. We add one to account for the spike that occurred before $\theta^{1}_{-2}$.
We plot contours as the number of estimated spikes transitions through the integers (Figure 11). Observe that the estimated boundaries follow the shape of the actual count boundaries and remain in proximity to these boundaries. The error in estimation is attributed to the limitations of this approach outlined in the beginning of this section, namely that the estimates are based off of quantities derived in the $\varepsilon\to 0$ limit.

### v2

@@ 3.4 Deriving the super-critical manifold
$\mathcal{Z}=\left\{(u,v,\theta_{0})\in\mathbb{R}^{2}\times[0,2\pi)\,|\,v=F(u),\ u-bv+\mu-Eb\cos(\theta_{0}-\pi/2)=0\right\}.$ (22)
@@
Figure 4: Bursting solution with critical and supercritical manifold. A The critical manifold, $\mathcal{M}$, is separated into attracting ($\mathcal{M}_{a-}$, $\mathcal{M}_{a+}$; blue) and repelling ($\mathcal{M}_{r}$; red) regions. The attracting and repelling regions intersect along fold lines, $\mathcal{L}_{\pm}$. The folded node ($N\ell$) and folded saddle ($S\ell$) are shown as a square and circle, respectively, along the fold lines. The supercritical manifold, $\mathcal{Z}$ (magenta), is a curve along the critical manifold. A bursting trajectory of system (1) with three spikes (forcing parameters $E=0.55$, $\omega=0.0149354$) is shown with $\theta=\omega t$ over one input period. This trajectory contains canard segments, following passage near the folded node and folded saddle, that remain near $\mathcal{M}_{a+}$ for non-negligible time on the fast timescale. B The same trajectory projected onto the $\theta$$x$ axis. Colors and symbols follow the same conventions as in panel A.
@@
00 0 $=-v+F(u)$ (23a)
00 0 $=u-bv+\mu-Eb\cos\left(\theta-\pi/2\right)$ (23b)
$\frac{d\theta}{d\tau_{2}}$ $=1$ (23c)
@@
$J_{SL}=\begin{bmatrix}\varepsilon^{-1}F^{\prime}(u)&-\varepsilon^{-1}\\ 1&-b\end{bmatrix}$ (24)
@@
Converting to the original $(x,y,\theta)$ coordinates (with $x=u-1$, see Sec. 3.1), a Hopf bifurcation occurs at $x_{dh,\pm}=\pm\sqrt{1-\varepsilon b}$ along $\mathcal{Z}$. This Hopf bifurcation does not belong to the original system (2), but instead to the slow layer problem (21). Such bifurcations are called delayed Hopf bifurcations [49, 50, 51, 24].
Remark 4
@@ 4 Spike-adding mechanisms in the FHN model with periodic forcing
[…]
First, we compare our theoretical results (summarized in Figure 3) with our numerical observations of the FHN model with periodic input (shown in Figure 2). We find that parametric regions where bursting solutions maintain a given number of spikes, and the boundaries between them, appear primarily between the curves of folded saddle-nodes on $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$ respectively; see Figure 5. This area is the union of regions II and III where the model has two and only two folded singularities (folded saddle plus folded node/focus), with both of them located on $\mathcal{L}_{-}$.
[…]
@@
To conclude, we will work in the remaining sections of the paper (Sec. 4.3 and beyond) under the assumption that $E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}$ see Figure 5, the region between the two curves labeled “folded SN on $\mathcal{L}_{-}$” and “folded SN on $\mathcal{L}_{+}$.
Figure 5: Parameter region of interest for bursting. Cusp-like boundaries for changing spike counts are located where a folded node and folded saddle exist on $\mathcal{L}_{-}$ (region II in Figure 3; see also Figure 2). The regions below the saddle-node (SN) bifurcation on $\mathcal{L}_{-}$ ($E=E^{*}_{\ell,\delta}$) and above the curve marking the transition from folded node to folded focus ($E=E^{**}_{\ell,\delta}$) are grayed-out. The parameter region above the SN on $\mathcal{L}_{+}$ ($E=E^{*}_{r,\delta}$) is shown in white. In our analysis we constrained input forcing parameters to $E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}$. Note that the horizontal axis closely resembles that of Figure 2 where $\omega=\varepsilon\delta\in[0,0.08]$ and $\varepsilon=0.08$.
@@ 4.2 Complementary FHN dynamics in other input-based regions of the parameter plane
[…]
As also shown in Figures 1 and 2 numerically, Region I corresponds to parameter values for which the system does not admit bursting solutions. Therefore, we do not consider this region in our study.
@@ 4.4 Adding spikes through the folded saddle canard
By Proposition 1, the desingularized system (10) admits a folded saddle on $\mathcal{L}_{-}$ as long as $E>E^{*}_{\ell,\delta}$. However, the desingularized system was obtained by a local state-dependent time rescaling that reverses the direction of the flow on $\mathcal{M}_{r}$ compared to the flow of (9) and (1). As such, desingularized solutions beginning in $\mathcal{M}_{r}$ that flow towards the folded saddle point $P_{S\ell}$ of coordinates $(u_{S\ell},v_{S\ell},\theta_{S\ell})$ (see Proposition 1) are reversed in the slow-reduced and in the full problem. In particular, the points that lie on $W_{S\ell}^{s}\cap\mathcal{M}_{r}$, where $W_{S\ell}^{s}$ is the stable manifold of the folded saddle, are repelled in the full system (1) away from $P_{S\ell}$. The curve $W_{S\ell}^{s}$ provides a path for slow solutions to pass through the folded saddle from $\mathcal{M}_{a-}$ into $\mathcal{M}_{r}$. This curve forms a vrai canard (or true canard), which perturbs to a canard segment in $\mathcal{M}_{a-}^{\varepsilon}\cap\mathcal{M}_{r}^{\varepsilon}$ for $\varepsilon>0$ [36].
For the desingularized system (10) the vrai canard acts as a separatrix, distinguishing solutions that pass between $\mathcal{M}_{a-}$ and $\mathcal{M}_{r}$ from those that do not. For the full system (1) the canard serves as a threshold, distinguishing solutions that remain in the neighborhood of $P_{S\ell}$ while crossing into $\mathcal{M}_{r}^{\varepsilon}$ from those that jump across to reach $\mathcal{M}_{a+}^{\varepsilon}$ immediately [53, 38].
Figure 7: Spike adding at folded saddle. Trajectories from the full three-dimensional system (3) projected onto the $\theta$ – $x$ plane for $E=0.482$ and (A) $\omega=0.02206875$, (B) $\omega=0.0220625$, and (C) $\omega=0.02$ (all in Region II of Figure 3). As a trajectory nears the termination of a burst sequence, its intersection with $\mathcal{M}_{a-}$ approaches $W^{s}_{S\ell}$. The proximity to $W^{s}_{S\ell}$ determines if there is a jump-back canard (A, no spike), a jump across canard (B, spike), or a jump at the fold curve (C, spike) followed by a subsequent return to the super-slow manifold $Z$ (in purple). The transition from jump-back to jump-across canard marks the addition of a spike near the folded saddle (indicated by a circle). In panel B (resp. C) the trajectory from A (resp. B) is shown in gray for comparison. In the desingularized system, flows are reversed along $\mathcal{M}_{r}$; the stable manifold of the folded saddle $W^{s}_{S\ell}$ (in blue) and the unstable manifold of the folded saddle $W^{u}_{S\ell}$ (in red; very close to $Z$) are drawn as dashed lines in this region.
@@
If no member of the $\theta$-sequence is close to the intersection of the separatrix with $\Sigma_{S\ell,-2}$, then we do not consider spike-adding to occur at the folded saddle. In this case, upon reaching $\theta^{m-1}_{-2}$, a solution flows towards the fold line $\mathcal{L}_{-}$ and makes a fast jump (relaxation oscillation) at it. The return dynamics to $\mathcal{M}_{a-}^{\varepsilon}$ bring the solution well past separatrix, leading it back to $\mathcal{Z}^{\delta}$, as in the previous cases (Figure 7C).
@@ 4.5 Adding spikes through the folded node canard
At the onset of an input period, bursting solutions are funneled towards the repelling region of the slow manifold through the folded node $P_{N\ell}$ of coordinates $(u_{N\ell},v_{N\ell},\theta_{N\ell})$ (see Proposition 1). Upon exiting the folded node, trajectories remain near the repelling manifold, forming canard segments at the start of a burst. The segment escapes the repelling region by either jumping back to the lower attracting branch where it originated (Figure 8A, B; see the trajectory immediately past the square mark), or by jumping across to the upper attracting branch (Figure 8C):
@@
A jump-back out of the folded node does not necessarily decrease the total number of spikes possible in a bursting oscillation. Suppose, for example, that an initial jump-across occurs and $n-1$ spikes follow (giving a total of $n$ spikes) before spike-termination at the folded saddle. A fast jump-back and subsequent return to the fold curve may leave enough time (or phase $\theta(t)$) for $n$ spikes before termination at the folded saddle. To alter the number of spikes in a burst, the initial jump-back must significantly change the trajectory’s ultimate proximity to $W_{S\ell}^{s}$, the stable manifold of the folded saddle, to undergo the spike-changes described in the previous section.
@@
Figure 8: Spike adding at folded node. Trajectories from the full three-dimensional system (3) projected onto the $\theta$ – $x$ plane for $E=0.482$ and (A) $\omega=0.0236$, (B) $\omega=0.02506875$, and (C) $\omega=0.025075$ (all in Region II of Figure 3). Trajectories are funneled towards the folded node (square) and into $\mathcal{M}_{r}$. After an excursion along $\mathcal{M}_{r}$, trajectories either jump back towards $\mathcal{M}_{a-}$ (A, B, no spike) or jump across towards $\mathcal{M}_{a+}$ (C, spike). From A to B, the escape from $\mathcal{M}_{r}$ towards $\mathcal{M}_{a-}$ yields a small amplitude oscillation with increasing magnitude. From B to C, the small amplitude oscillation gives rise to a jump-across towards $\mathcal{M}_{a+}$. The jump-across marks the addition of a spike out of the folded node canard. In panel B (resp. C) the trajectory from A (resp. B) is shown in gray for comparison. In the desingularized system, flows are reversed along $\mathcal{M}_{r}$, so the stable manifold of the folded saddle (circle) which is $W^{s}_{S\ell}$ (in blue), and the unstable manifold $W^{u}_{S\ell}$ of the folded saddle (in red; very close to the super-critical manifold $Z$) are drawn as dashed lines in this region.
Remark 5
Folded node canard dynamics occur exclusively in Region II of Figure 3. As the parameter regime transitions from Region II into III, the folded node becomes a folded focus and the canards exist only at the folded saddle. The cusps in the spike-count boundaries appear only in Region II as well (see Figure 5) -- they arise from the complex interaction of the jump-back dynamics at the folded node with the jump-across dynamics at the folded saddle. In Region III where there is a folded focus, the spike-adding is influenced by the decay rate of the stable spirals near the folded focus and the folded saddle-canard dynamics. However, these interactions do not induce cusp-like boundaries (transparent-blue area in Figure 5). In the next section, we examine the interaction between the folded node and folded saddle dynamics in more detail.
@@ 5.1 Level sets of trajectory $L^{2}$ norm
$\left(\frac{1}{T}\int_{0}^{\top}\left[x(t)^{2}+y(t)^{2}\right]\,\mathrm{d}t\right)^{1/2}$ (25)
@@
These two organizing features -- the asymptotic curves and the cusp-connecting level sets -- depict how the cusps form in the spike-count boundaries. Fundamentally, cusps occur in parameter regimes in which the bifurcation between the jump-back and jump-across canards at the folded node leads to trajectories that ultimately land close enough to the stable manifold of the folded saddle to be influenced by saddle-canard dynamics. Put another way, the cusps arise as the bifurcation curve for spike-adding at the folded node (Figure 9; see the curves depicted by the blue arrows) collides with the bifurcation curve for spike-adding at the folded saddle (Figure 9; see the curves depicted by red arrows). Thus, we attribute the cusp formation to the interaction of the folded node and folded saddle dynamics.
@@ 5.2 Phase distance
The mechanisms of spike-adding at the folded node and folded saddle indicate how the spike-count boundaries might be derived analytically. Estimating the time spent (or ‘phase’ spent when considering the $\theta$ – $x$ plane) on $\mathcal{M}_{a-}^{\varepsilon}$ before the first jump on the fast timescale gives an approximation for the delay-to-spike within a burst. Similarly, estimating where the stable manifold of the folded saddle intersects the line $x=-2$ allows an approximation for spike termination (recall that $\Sigma_{S\ell,-2}$ denoted the plane containing this point). In particular, one would consider the phase difference between these two estimates, relative to the intra-burst spiking frequency, to estimate the number of spikes in a burst.
Figure 10: Estimating duration of spiking phase. A sample trajectory’s return to the line $x=-2$ following its first spike is marked as $\theta^{1}_{-2}$. The intersection of the folded saddle’s stable manifold with $x=-2$ is marked as $\theta_{S\ell,-2}$. The difference between these values, $\Delta\theta$, was used to estimate the number of spikes in a burst; see Eq. (26). Circle/square markers and the color code for different curves are the same as in previous figures.
This approach is hindered by a few factors. First, predicting the bifurcation point between a jump-back and a jump-across canard at the folded node is difficult, as these dynamics likely involve bifurcations of canard solutions [33, 37]. Second, the stable manifold of the folded saddle comes from the singular limit as $\varepsilon\to 0$, making it hard to exactly predict the separatrix that distinguishes jump-back from jump-across solutions at the folded saddle. And third, we do not establish an analytic relationship between the intra-burst spiking frequency and input parameters.
@@
number of $\text{number of spikes}=1+\left[\frac{\Delta\theta}{1000\times\omega}\right]\times f_{burst}\,.$ (26)
For numerical calculations, we set $f_{burst}=27\,\mathrm{Hz}$ as this was the predominant frequency observed in the constantly forced system. Within formula (26), we round to the next largest integer as a fractional quantity still allows for an additional spike at the folded saddle. We add one to account for the spike that occurred before $\theta^{1}_{-2}$.
We plot contours as the number of estimated spikes transitions through the integers (Figure 11). Observe that the estimated boundaries follow the shape of the actual count boundaries and remain in proximity to these boundaries. The error in estimation is attributed to the limitations of this approach outlined in the beginning of this section, namely that the estimates are based off of quantities derived in the $\varepsilon\to 0$ limit. Indeed, when we repeated the relevant computations using smaller values of $\varepsilon$ (not shown), we observed that the discrepancy decreased accordingly.
@@
6 Periodic forcing in the Morris-Lecar model
Figure 12: Spike-adding in the Morris-Lecar model. A Spike-adding transitions as forcing frequency $f$ varies from 0.5 to 10 Hz and forcing amplitude from 10 to 70 with $E_{0}$ fixed at 80 (see Eq. (28)). Cusp-like transitions emerge as in the forced FHN model. The white box highlights the region with a non-monotonic relationship between the number of spikes and either forcing parameter. We omit the portion of the gray region (1-spike or no-spikes) up to $\omega=\phi$, or $f\approx 12\,\text{Hz}$. B A sample bursting trajectory (see white scatter point in zoomed in panel of A) containing putative canard-with-head segments near spike initiation and termination. C Bifurcation diagram of the unforced Morris–Lecar model, showing two subcritical Hopf bifurcations and nearby saddle-nodes of limit cycles, similar to the unforced FHN model (see Figure 1 in [29]).
The FHN model is a simplified dynamical model for neural dynamics which it is not based on obvious underlying biophysical principles. Nevertheless, when considering a more realistic model for action potential production – the Morris-Lecar (ML) model – we observed similar spike-adding transitions induced by changes in the input amplitude and/or frequency; see Figure 12.
The full equations for the ML model are shown in Appendix 8.4. Briefly, we follow the model presented in [22, Chapter 3], which is a two-dimensional model describing the dynamics for voltage, $V$, and potassium activation, $n$:
$C_{M}\dfrac{dV}{dt}$ $=I_{\text{app}}(t)-g_{\text{L}}(V-E_{\text{L}})-g_{\text{K}}n(V-E_{\text{K}})-g_{\text{Ca}}m_{\infty}(V)(V-E_{\text{Ca}})$ (27a)
$\dfrac{dn}{dt}$ $=\phi\left(n_{\infty}(V)-n\right)/\tau_{n}(V)$ (27b)
We represent the applied current, $I_{\text{app}}$ as a function of time but choose system parameters so that the ML system with constant forcing ($I_{\text{app}}(t)\equiv E_{0})$ exhibits subcritical Hopf bifurcations with nearby saddle-nodes of limit cycles; see Figure 12C. This bifurcation structure closely matches the structure of the studied FHN equations with constant input [29]. In the case of the ML model, the first subcritical Hopf point occurs for external input $E_{0}\approx 100$; for the FHN model the first Hopf point occurs when the external forcing magnitude is around 0.3. Therefore, we augment the periodic external current by shifting its baseline closer to the Hopf point for constant input,
$I_{\text{app}}=E_{0}+E\sin(\omega t).$ (28)
[…]
Like the FHN model with periodic forcing, solutions from the ML system with periodic forcing exhibit bursting patterns. Furthermore, some solutions contained small-amplitude oscillations near spike-initiation and termination, characteristics of the canard segments studied in the periodically forced FHN system (see Figure 12B). In simulations that swept across forcing parameters in ranges analogous to those used in the FHN simulations, similar spike-adding boundaries emerged. In particular, the cusp-like boundaries were present when $\omega=\mathcal{O}(\phi\sqrt{\phi})$ with the pattern terminating near $\omega=\phi^{2}$ in parameter space, similar to the $\mathcal{O}(\varepsilon\sqrt{\varepsilon})$ to $\mathcal{O}(\varepsilon^{2})$ studied in the FHN system. We do not perform the same detailed analysis of folded equilibria as in the FHN prototype, but we expect similar dynamics studied in the simpler FHN system to hold for the more biologically realistic Morris-Lecar model.